OpenStudy (tmagloire1):

If y = 3x3 – 2x and dx/dt equals 3, find dy/dt when x = –2. Give only the numerical answer. For example, if dy, dt = 4, type only 4.

OpenStudy (anonymous):

so y = 9 - 2x?

OpenStudy (tmagloire1):

how did you get that

OpenStudy (anonymous):

Sorry, I'm asking you to clarify, what do you mean by "y = 3x3 – 2x" is that 3 multiplied by 3 or is that 3 times x to the power 3?

OpenStudy (tmagloire1):

3x^3-2x

OpenStudy (tmagloire1):

sorry my fault

OpenStudy (anonymous):

no problem, well dy/dt = (dy/dx)*(dx/dt) (chain rule) then dy/dx = 9 x^2 - 2 we already know dx/dt = 3 therefore dy/dt = 3*(9x^2 -2) so dy/dt = 27x^2 - 6

OpenStudy (anonymous):

oh and then x = -2 so you get 102 i think

OpenStudy (tmagloire1):

Okay i had to implicitly differentiate i forgot. Thank you so much!

OpenStudy (anonymous):

np